OpenStudy (anonymous):

ingrate 2/(5x^2+4)...please draw thanks

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 2 }{ 5x ^{2}+4 }dx \rightarrow 2\int\limits_{}^{}\frac{ dx }{ ((\sqrt{5} x)^{2}+(2)^{2})} \rightarrow u=\sqrt{5}x,du=\sqrt{5} \rightarrow \frac{ 2 }{ \sqrt{5} } \int\limits_{}^{}\frac{ du }{ (\sqrt{5}x)^{2}+(2)^{2} }\]